Swimming pool design. An architect has designed a motel pool within a rectangular area that is fenced on three sides. If she uses 60 yards of fencing to enclose an area of 352 square yards, then what are the dimensions of L and W? Assume L is greater than W.
Answer by checkley77(12844)

L*W=352 OR L=352/W
L+2W=60
352/W+2W=60
(352+2W*W)/W=60
(352+2W^2)=60W
2W^2-60W+352=0
2(W^2-30W+176)=0
2(W-22)(W-8)=0
W-8=0
W=8 FOR THE WIDTH.
L+2*8=60
L+16=60
L=60-16
L=44 FOR THE LENGTH.
PROOF
44*8=352
352=352